OpenStudy (anonymous):

through(1,-5) , slope (2,-3)

hero (hero):

Using y = mx + b, insert points (1,-5) and (2,-3) to get: -5 = 1m + b -3 = 2m + b Now, isolate b: b = -5 - 1m b = -3 - 2m Now, set expression b = b -5 - 1m = -3 - 2m Now, place like terms on the same side using inverse operations: 2m - m = 5 - 3 m = 2 Now find b: b = -5 - 1m b = -5 - 1(2) b = -5 - 2 b = -7 Therefore the equation of the line is y = 2x - 7

OpenStudy (anonymous):

can you help me with one more ?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

I thought I had done this one already.

hero (hero):

@bubblekush, I showed you the steps. You should be able to figure out these kinds of problems yourself now. All you do is insert the points into y = mx + b, then follow the steps I posted.